Which fo the following are isoelectronic and isostructural?
NO3-, CO32 -, CIO3-, SO3
a. NO3-, CO32 -,
b. SO3, NO3-
c. CIO3-, CO32 –
d. CO32 -, SO3

Grade:upto college level

4 Answers

Deepak Patra
9 years ago

(a) NOTE : Isoelectronic species have same number of electrons and isostructural species have same type of hybridization at central atom.

NO3; No. of e = 7 + 8 * 3 + 1 = 32, hybridization of N in NO3 is sp3

CO32–; No. of e = 6 + 8 * 3 + 2 = 32, hybridization of C in CO32– is sp3

CIO3; No. of e = 17 + 8 * 3 + 1 = 42, hybridization of CI in CIO3 is sp3

SO3; No. of e = 16 + 8 * 3 = 440, hybridization of S in SO3 is sp2

? NO3 and CO32– are isostructural and isoelectronic.

Arun Kumar IIT Delhi
9 years ago
Hi
Its A.
Both have 32 e-.
Both have trigonal planar structure.
Thanks & Regards, Arun Kumar, Btech,IIT Delhi, Askiitians Faculty
avanish
6 years ago
Correct answer is ABecause both molecule have same number of total electrons and same hybridisation sp3
Yash Chourasiya
3 years ago
Dear Student

NO3–; No. of e–= 7 + 8 * 3 + 1 = 32,
Hybridization of N in NO3–is sp3
CO32–; No. of e–= 6 + 8 * 3 + 2 = 32,
Hybridization of C in CO32–is sp3
CIO3–; No. of e–= 17 + 8 * 3 + 1 = 42,
Hybridization of CI in CIO3–is sp3
SO3; No. of e–= 16 + 8 * 3 = 440,
Hybridization of S in SO3is sp2

Hence option A is correct.

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya